# Reconfigurable Intelligent Surfaces for 6G Mobile Networks: An Industry R&D Perspective

### v2

@@ IV-B Deployment Optimization
Developing an RIS-assisted wireless communication networks raises the problem of RIS placement, a part of the so-called network planning problem. In order to realize RIS placement, a model that captures the main propagation characteristics of RIS is needed, while still being mathematically tractable given the large scale of network planning problems. The question of “what level of abstraction” is then raised in order to maintain the right balance between accuracy and complexity.
@@
Migration to higher frequency bands such as millimeter wave (mmWave) or the sub-terahertz (sub-THz) to achieve the ambitious requirements in 6G wireless networks on the one hand, and the product-distance path loss model of RIS on the other hand, imply the need for large surfaces with many elements, which can achieve passive beamforming gain through joint reflection. To compensate for the severe path loss, besides the required large surface, the existence of the LOS path is indispensable, which is equivalent to the LOS link between BS-RIS and RIS-UE. In addition to improving the channel gain, the LOS link also simplifies the control and operation of an RIS, which is a typical assumption in the literature. This issue, moreover, explains a prospective use case of RISs, namely coverage extension, by providing virtual LOS links through the RISs.
@@
Furthermore, another important feature that is largely neglected in the literature is the orientation of the surface. More specifically, the orientation of the surface with respect to the BS and UE. It is known that the further we move away from the surface broadside (i.e., toward the end fire), the higher the effects of impairments such as beam squint and gain loss become. Therefore, in a realistic model, we need to limit the field of view (FoV) of the surface, which is determined, for instance, based on the power budget and ensures that the BS and considered area of interest (i.e., technically where UEs are supposed to be) are within this FoV.
@@
We consider a multi-RIS-assisted wireless communication system in a scenario as shown in Figure 8, which comprises a BS located at $P_{B}=[b_{x},b_{y},b_{z}]^{T}$, an RIS deployment area $P=\{P_{1},\ldots,P_{M}\}$ with $M$ RIS candidate locations, and an area of interest $L=\{L_{1},\ldots,L_{N}\}$ with $N$ potential locations for UEs, to be covered either by the BS or an RIS. We assume a LOS link to the BS location, $P_{B}$, for all $P\in\mathcal{P}$. Moreover, regarding the orientation of the surface, we define the valid FoV of the surface as $[-\theta_{\max},\theta_{\max}]$ with $\theta_{\max}$ being the maximum acceptable deviation from the broadside (for both incident and reflecting angles), that satisfies the given system constraints (e.g., requirements on the maximum beam squint effect and gain loss).
@@
For each candidate location of the RIS, the range of valid orientations of the surface is determined in a way to ensure the steering angle towards the BS is in the range of valid FoVs. Figure 9 illustrates an example of a valid range of azimuth angles for an arbitrary candidate position of RIS and the given BS location. The blue surface (line in the figure) is the one with broadside in the direction of the BS, and the red and green ones are the ones that reach ±$\pm\theta_{\max}$ in steering angle toward the BS. Therefore, the valid range of orientations of the surface that provide a steering angle within the valid FoV are all orientations between the orientation of red and green surfaces. Similarly, for each candidate location of the RIS and for a given UE position, the range of valid orientations of the surface is calculated. This procedure should be repeated for the whole UE positions that are supposed to be covered by the considered RIS location (i.e., a subset of $L$). As a result, for each RIS candidate location, given the BS location, and the set of UE positions that are supposed to be covered by the considered RIS, the final valid orientations of the surface are calculated by the intersection of all individually determined valid orientations of the surface, i.e., given by BS and the subset UE positions.
@@
In the following, we illustrate a toy example of extended coverage provided by RISs demonstrating the importance of placement and orientation of the surfaces. For the purpose of the toy example, we use the channel model QuaDRiGa [29, 30] developed at Fraunhofer HHI, which has recently been extended by the feature of LOS-detection. This allows site-specific simulations by incorporating data from the environment. Here, the data includes 3D-information of buildings and height profiles of the terrain under consideration, which is the campus network “Berlin 5G testbed”, shown in Figure 10. The location of the BS, $P_{B}$ is on top of the HHI building in Berlin, Germany.
@@
Here, we consider a small-scale scenario in which a set $\mathcal{P}=\{P_{1},\ldots,P_{5}\}$ of 5 RIS candidate positions, each at a height of 30 meters, is available. Their locations and respective coverage are shown in Figure 12. White bars represent the RIS (not the right scale with respect to the buildings) and the respective total colored areas indicate the LOS coverage by the corresponding RIS without taking into account the RIS orientation. Darker colored areas show the LOS coverage with the RIS orientation that maximizes LOS coverage for and dashed white lines correspond to the respective broadside of the deployed RIS.
Figure 12: LOS coverage of IRS candidate locations.
@@
The deployment for the optimal RIS, i.e., the one that extends the coverage the most, is shown in Figure 13. In Figure 13 (bottom), the second RIS is deployed which optimally (in a greedy sense) extends the superposition of the LOS coverage areas of the BS and the first RIS.
@@ IV-C Dimensioning
An efficient deployment strategy of RISs in mobile networks includes not only the selection of the suitable locations and orientations of the RISs but also their dimensioning because the number of elements and the size of the RISs significantly determine the achievable reflection coefficient. Initial modeling of path loss can provide information for dimensioning (and be extended in more comprehensive link budget analyses). For this purpose, consider the scenario of an RIS-supported network from Figure 14, which includes an RIS with $N$ elements and size $L\times L\,\text{m}^{2}$. The distances of the three paths $\rho_{d}$, $\rho_{t}$, and $\rho_{r}$, which correspond to the path between transmitter and receiver, transmitter and RIS, and RIS and receiver, respectively, each has a free-space path loss (FSPL) exponent of 2. In addition, the direct path between the transmitter and receiver is blocked, so the signal experiences attenuation at a certain level.
In Figure 15 the resulting path losses are plotted against the distance $\rho_{r}$ for $\rho_{t}=20\,\text{m}$ in (a) and for $\rho_{t}=\rho_{r}$ in (b) at a frequency of 6 GHz. While an RIS with $N=121$ and $L=0.275$ is not sufficient to compensate for the path loss of the direct path with a $20$ dB blockage, this is possible with an RIS with $N=484$ and $L=0.55\,\text{m}$. However, this RIS configuration also lags behind a direct LOS link without blockage as well as behind a LOS link with 20 dB blockage at larger distances $\rho_{t}=\rho_{r}\geq 85\,\text{m}$.
(a) Number of elements
(b) RIS length
@@
elements, with an assumed element size of $\lambda/2$, is required, so that the free-space path loss of the RIS-assisted link is equal to that of the LOS link without blockage.
Based on this, the required number of elements and required RIS size is shown in Figure 16 (over distance $\rho_{r}=\rho_{t}$ and for a frequency of 6 GHz). It can be seen that compensation for path loss, especially for the non-blocked LOS path or high distances, requires a large number of elements or a very large RIS. This is primarily because the RIS-aided link suffers from double path loss, which is accordingly referred to as product-distance path loss [32].
@@
Figures 17 and 18 correspond to Figures 15 and 16, respectively, but for a frequency of 28 GHz. The trends are generally consistent with the 6 GHz calculations. In addition, by comparison, it can be seen that the signals generally experience greater attenuation at higher frequencies. Due to the product-distance path loss, far more elements are then required (625 elements are not sufficient here to compensate for the path loss of the LOS link with 20 dB blockage, but 2500 elements are) than at a frequency of 6 GHz. At the same time, this can be achieved with smaller RISs (approximately $L=0.26786$ m), since the size of the elements and thus the overall size of the RIS is frequency dependent, e.g., assuming an element size of $\lambda/2$, where $\lambda$ corresponds to the RIS wavelength of operation.
(a) Number of elements
(b) RIS length
@@
V Conclusions and Outlook
This paper presents an overview of RIS technology from and Industry R&D perspective under the scope of the 6G LICRIS project. Potential use cases, technology requirements, and existing challenges are outlined. RIS hardware concepts are presented and practical problems such as RIS placement and dimensioning as well as radio propagation modeling are discussed.
@@ V Conclusions and Outlook
Overall, while RIS is a promising technology for 6G, various practical problems need to be addressed before incorporating such a solution into mobile cellular systems. In particular, RIS channel estimation, control signalling overhead, and coexistence concerns are some of the serious challenges that remain largely unsolved.

### v4

@@ IV-B Deployment Optimization
Developing an RIS-assisted wireless communication network raises the problem of RIS placement, a part of the so-called network planning problem. In order to realize RIS placement, a model that captures the main propagation characteristics of RIS is needed, while still being mathematically tractable given the large scale of network planning problems. The question of “what level of abstraction” is then raised in order to maintain the right balance between accuracy and complexity.
@@
Migration to higher frequency bands such as millimeter wave (mmWave) or the sub-terahertz (sub-THz) to achieve the ambitious requirements in future wireless networks on the one hand, and the product-distance path loss model of RIS on the other hand, imply the need for large surfaces with many elements, which can achieve passive beamforming gain through joint reflection. To compensate for the severe path loss, besides the required large surface, the existence of the LOS path is indispensable, which is equivalent to the LOS link between BS–RIS and RIS–UE. In addition to improving the channel gain, the LOS link also simplifies the control and operation of an RIS, which is a typical assumption in the literature. This issue, moreover, explains a prospective use case of RISs, namely coverage extension, by providing virtual LOS links through the RISs.
@@
Furthermore, another important feature that is largely neglected in the literature is the orientation of the surface. More specifically, the orientation of the surface with respect to the BS and UE. It is known that the further one moves away from the surface broadside (i.e., toward the end fire), the higher the effects of impairments such as beam squint and gain loss become. Therefore, in a realistic model, the field of view (FoV) of the surface needs to be limited, which is determined, for instance, based on the power budget and ensures that the BS and considered area of interest (i.e., technically where UEs are supposed to be) are within this FoV.
@@
A multi-RIS-assisted wireless communication system is considered in a scenario as shown in Figure 11, which comprises a BS located at $\textbf{P}_{B}=[b_{x},b_{y},b_{z}]^{T}$, an RIS deployment area $P=\{P_{1},\ldots,P_{M}\}$ with $M$ RIS candidate locations, and an area of interest $L=\{L_{1},\ldots,L_{N}\}$ with $N$ potential locations for UEs, to be covered either by the BS or an RIS. A LOS link to the BS location, $\textbf{P}_{B}$, for all $P\in\mathcal{P}$ is assumed. Moreover, regarding the orientation of the surface, the valid FoV of the surface is defined as $[-\theta_{\max},\theta_{\max}]$ with $\theta_{\max}$ being the maximum acceptable deviation from the broadside (for both incident and reflecting angles), that satisfies the given system constraints (e.g., requirements on the maximum beam squint effect and gain loss).
@@
For each candidate location of the RIS, the range of valid orientations of the surface is determined in a way to ensure the steering angle towards the BS is in the range of valid FoVs. Figure 12 illustrates an example of a valid range of azimuth angles for an arbitrary candidate position of RIS and the given BS location. The blue surface (line in the figure) is the one with broadside in the direction of the BS, and the red and green ones are the ones that reach ±$\pm\theta_{\max}$ in steering angle toward the BS. Therefore, the valid range of orientations of the surface that provide a steering angle within the valid FoV are all orientations between the orientation of red and green surfaces. Similarly, for each candidate location of the RIS and for a given UE position, the range of valid orientations of the surface is calculated. This procedure should be repeated for all UE positions that are supposed to be covered by the considered RIS location (i.e., a subset of $L$). As a result, for each RIS candidate location, given the BS location, and the set of UE positions that are supposed to be covered by the considered RIS, the final valid orientations of the surface are calculated by the intersection of all individually determined valid orientations of the surface, i.e., given by BS and the subset UE positions.
@@
In the following, a toy example of extended coverage provided by RISs is illustrated, demonstrating the importance of placement and orientation of the surfaces. For the purpose of the toy example, the channel model “Quasi Deterministic Radio Channel Generator” QuaDRiGa [49, 50] developed at Fraunhofer Institute for Telecommunications, Heinrich-Hertz-Institut (HHI) is used, which has recently been extended by the feature of LOS-detection. This allows site-specific simulations by incorporating data from the environment. Here, the data includes 3D information of buildings and height profiles of the terrain under consideration, which is the campus network “Berlin 5G testbed”, shown in Figure 13. The location of the BS, $P_{B}$, is on top of the HHI building in Berlin, Germany.
@@
Here, a small-scale scenario is considered in which a set $\mathcal{P}={P_{1},\ldots,P_{5}}$ of $5$ RIS candidate positions, each at a height of $30$ meters, is available. Their locations and respective coverage are shown in Figure 15. White bars represent the RIS (not to the right scale with respect to the buildings) and the respective total colored areas indicate the LOS coverage by the corresponding RIS without taking into account the RIS orientation. Darker colored areas show the LOS coverage with the RIS orientation that maximizes LOS coverage, and dashed white lines correspond to the respective broadside of the deployed RIS. The deployment for the optimal RIS, i.e., the one that extends the coverage the most, is shown in Figure 16. In the bottom part of this figure, the second RIS is deployed, which optimally (in a greedy sense) extends the superposition of the LOS coverage areas of the BS and the first RIS.
Figure 15: LOS coverage of RIS candidate locations.
@@ IV-C Dimensioning
An efficient deployment strategy for RISs in mobile networks includes not only the selection of suitable locations and orientations for the RISs but also their dimensioning, because the number of elements and the size of the RISs significantly determine the achievable reflection coefficient. Initial modeling of path loss can provide information for dimensioning (and be extended in more comprehensive link budget analyses). For this purpose, consider the scenario of an RIS-supported network from Figure 17, which includes an RIS with $N$ elements and size $L\times L\,\text{m}^{2}$. The distances of the three paths $\rho_{d}$, $\rho_{t}$, and $\rho_{r}$, which correspond to the path between the transmitter and receiver, the transmitter and RIS, and the RIS and receiver, respectively, each have a free-space path loss (FSPL) exponent of 2. In addition, the direct path between the transmitter and receiver is blocked, so the signal experiences attenuation at a certain level.
In Figure 18, the resulting path losses are plotted against the distance $\rho_{r}$ for $\rho_{t}=20\,\text{m}$ in (a) and for $\rho_{t}=\rho_{r}$ in (b) at a frequency of $6$ GHz. While an RIS with $N=121$ and $L=0.275\,\text{m}$ is not sufficient to compensate for the path loss of the direct path with a $20$ dB blockage, this is possible with an RIS with $N=484$ and $L=0.55\,\text{m}$. However, this RIS configuration also lags behind a direct LOS link without blockage as well as behind a LOS link with $20$ dB blockage at larger distances $\rho_{t}=\rho_{r}\geq 85\,\text{m}$.
(a) Number of elements for $\rho_{t}=20\,\text{m}$
(b) RIS length for $\rho_{t}=\rho_{r}$
@@
of elements, with an assumed element size of $\lambda/2$, is required so that the free-space path loss of the RIS-assisted link is equal to that of the LOS link without blockage.
Based on this, the required number of elements and required RIS size are shown in Figure 19 (over distance $\rho_{r}=\rho_{t}$ and for a frequency of $6$ GHz). It can be seen that compensation for path loss, especially for the non-blocked LOS path or high distances, requires a large number of elements or a very large RIS. This is primarily because the RIS-aided link suffers from double path loss, which is accordingly referred to as product-distance path loss [52].
@@
Figures 20 and 21 correspond to Figures 18 and 19, respectively, but for a frequency of 28 GHz. The trends are generally consistent with the 6 GHz calculations. In addition, by comparison, it can be seen that the signals generally experience greater attenuation at higher frequencies. Due to the product-distance path loss, far more elements are then required ($625$ elements are not sufficient here to compensate for the path loss of the LOS link with 20 dB blockage, but $2500$ elements are) than at a frequency of 6 GHz. At the same time, this can be achieved with smaller RISs (approximately $L=0.26786$ m), since the size of the elements and thus the overall size of the RIS is frequency-dependent, e.g., assuming an element size of $\lambda/2$, where $\lambda$ corresponds to the RIS wavelength of operation.
(a) Number of elements for $\rho_{t}=20\,\text{m}$
(b) RIS length for $\rho_{t}=\rho_{r}$
@@
V Discussion
This section discusses the results presented in Section III about hardware and in Section IV about deployment. It is intentional that only the first steps and ideas of the 6G-LICRIS project are presented. The results will be presented in dedicated future publications regarding the hardware realization with LC, the propagation modeling, and system architecture aspects, as well as the deployment optimization, especially with the use and description of the channel model QuaDRiGa including RIS.
In Section III about the hardware, the RIS topology was presented where the realization was with liquid crystal. In most cases, RISs are realized with semiconductor-based approaches, commonly with PIN diodes or varactor diodes, and only a relatively few works exist with the realization of RIS with liquid crystal [53, 54, 55]. Compared to semiconductor-based RIS, LC is more energy-efficient in higher frequencies, it is scalable, cost-effective, and a promising technology candidate for RIS hardware implementation.
Two different types of RIS were presented: The first one is the multi-resonance RIS. The novelty of it is the large wideband, since most designs have a relative bandwidth of around 10%, the presented design boasts more than 20% bandwidth. It is also compatible with very thin LC-layers, which makes the tuning time of the LC much faster and reduces the amount of LC needed, and therefore, is capable of reducing costs.
Regarding the delay-line based RIS, it is comparable, for example, with Reference [54], but these approaches are based on one single polarization only. Theoretically, this approach should be compatible with dual polarization by placing two perpendicular lines. However, in order to obtain the required phase shift, the delay-lines need to be quite long, and it is not possible to place two lines in the same unit-cell. What is proposed is to significantly reduce the dimension that the delay-lines need to obtain the phase difference, and therefore, the required space to place two lines is achieved. Unfortunately, this also implies an increase in losses.
In Section IV-A, the propagation modeling was presented. This suits the integration of a RIS into a mobile network. A RIS integration mechanism with the existing 3GPP network architecture is proposed in the following. The approach for this is to use, for example, an open radio access network (O-RAN) architecture [56] to include the RIS orchestrator and controller. The key components within the O-RAN architecture are the Radio Unit (RU) domain and the Central Unit (CU) domain, with an additional intelligent controller (IC) function. In addition, the non-real-time radio access network intelligent controller (Non-RT RIC) logical function resides within the Service Management and Orchestration (SMO) framework and provides non-real-time optimization and control of RAN resources. It facilitates functionalities like artificial intelligence/machine learning (AI/ML) model training and policy-based management for near-real-time RIC applications. The RIS is now integrated regarding RIS orchestration in the Non-RT RIC and regarding RIS control in the near-real-time RIC. The advantage of this approach is that no new interfaces are needed. For the AI/ML model, approaches as are used. A detailed description of such models, as well as the architecture including RIS, is planned in future publications.
A deployment optimization was presented in Section IV-B as well. A channel model is used to calculate the coverage in an urban scenario for the system consisting of BS-UE and BS-RIS-UE compared to BS-NCR-UE [57]. The coverage is shown as a heatmap of the SNR. The findings suggest that a RIS is suitable for open spaces such as squares, while the NCR shows better improvements for narrow scenarios such as streets. In the present publication, only the RIS without NCR consideration for 2-dimensional scenarios is presented. Also, only the improved area is shown without a detailed signal strength or heatmap. However, the present publication demonstrated different RIS locations to optimize the deployment, which is missing in [57]. There are even deep reinforcement learning/artificial intelligence approaches for the deployment optimization, however, only for indoor applications [58]. Outdoor measurements of a RIS at $26$ GHz, constructed with PIN diodes and equipped with $1024$ elements and a predefined codebook, are presented in reference [8]. They provide a coverage analysis both with and without the RIS. The outdoor measurement only should be compared with the planned future publications.
The RIS dimensioning was presented in Section IV-C. This was an important consideration for the specification and the RIS size, which was also considered in Section III. One of the main findings is that a high number of RIS elements, $N$, is helpful to compensate for the low channel gain. This is in accordance with, for example, Reference [37] (see also Section II-C3).
Recently, new technologies that extend the RIS concept have emerged. For example, stacked intelligent metasurfaces adopt a different topology to perform computations in the wave domain, allowing for more degrees of freedom in signal processing [59, 60]. While the concept is conceptually promising, it might bring several engineering challenges that hinder its adoption in commercial products. For example, the stacked design incurs additional transmission losses that are not acceptable in real products. Furthermore, the meta-atoms might introduce non-linearity distortion in the system that complicates conformance to regulation and out-of-band emissions. In any case, these issues are open problems that remain to be solved in future work.
VI Conclusions and Outlook
This paper presents an overview of RIS technology from an industry R&D perspective under the scope of the 6G-LICRIS project. Potential use cases, technology requirements, and existing challenges are outlined. RIS hardware concepts are presented, and practical problems such as RIS placement and dimensioning, as well as radio propagation modeling, are discussed. A hardware design of a RIS with liquid crystal at 28 GHz is presented. A propagation model for RIS as a new part of the system architecture is discussed with the approaches of semi-empirical models, geometric models, and their combination through the application of artificial intelligence/machine learning. Finally, a channel model for deployment optimization and dimensioning is presented, with the findings that a rather large RIS is favored for coverage improvement as well as greater attenuation at higher frequencies combined with a smaller RIS size. The 6G-LICRIS project results will be presented in dedicated future publications regarding the hardware realization with LC, the propagation modeling, and system architecture aspects, as well as the deployment optimization, especially with the use and description of the channel model QuaDRiGa including RIS.
@@ VI Conclusions and Outlook
Overall, while RIS is a promising technology for 6G, various practical problems need to be addressed before incorporating such a solution into mobile cellular systems. In particular, RIS channel estimation, control signaling overhead, and coexistence concerns are some of the serious challenges that remain largely unsolved and shall be addressed in future work.